OpenStudy (anonymous):

Explain the distance formula. Then use it to calculate the distance between A(1, 1) and B(7, -7)

OpenStudy (michele_laino):

part B) the requested distance is: \[\Large \begin{gathered} d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( {7 - 1} \right)}^2} + {{\left( { - 7 - 1} \right)}^2}} = \sqrt {{6^2} + {{\left( { - 8} \right)}^2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

Part A)  If we apply the theorem of Pitagora, we can write: \[\Large d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} \]